# Semantically Diverse Language Generation for Uncertainty Estimation in Language Models

### v1

@@ 5 Related Work
Several works utilized the language model itself to obtain a prediction of their uncertainty, whether that be numerical or verbal (Mielke et al., 2022; Lin et al., 2022b; Kadavath et al., 2022; Cohen et al., 2023a; Ganguli et al., 2023; Ren et al., 2023; Tian et al., 2023). Cohen et al. (2023b) utilize cross-examination, where one language model generates the output sequence and the other language model acts as an examiner to assess the uncertainty. Zhou et al. (2023) investigate the behavior of language models when expressing their (un)certainty.
A large body of work focuses on sampling a set of output sequences to obtain sampling-based uncertainty estimators. Xiao and Wang (2021); Malinin and Gales (2021); Hou et al. (2023) incorporate both aleatoric and epistemic estimates of uncertainty, where epistemic uncertainty due to model selection is considered. While Kuhn et al. (2023); Lin et al. (2023); Duan et al. (2023) evaluate the aleatoric uncertainty only under a single given language model, they take the semantic equivalence of potential output sequences into account. Manakul et al. (2023) also sample a set of output sequences but utilize them as input to another language model to assess the uncertainty. Another approach to uncertainty estimation in NLG is conformal prediction (Quach et al., 2023), where a stopping rule for generating output sequences is calibrated. Additionally, Xiao et al. (2022) empirically analyze how factors such as model architecture and training details influence the uncertainty estimates in language models.
@@
Li et al. (2016) propose an alternative training procedure of language models to avoid generic, input-independent output sequences and increase diversity. Diverse beam search (Vijayakumar et al., 2018) optimizes for a diversity-augmented objective across beam groups, based on diversity heuristics. Ippolito et al. (2019) compare diversity encouraging decoding strategies. Nucleus sampling (Holtzman et al., 2020) generates higher quality as well as more diverse output sequences but does not explicitly encourage semantic diversity. Contrastive decoding (Li et al., 2023) utilizes a second, weaker language model, where the decoding algorithm favors tokens generated by the stronger model and penalizes tokens generated by the weaker model. Tam (2020) utilize semantic clustering during beam search, which is used to prune beams and diversify the remaining candidates. However, this only indirectly steers towards more diversity and relies on the diversity of the initial beams.
Closely related, but not directly targeting semantic diversity of output sequences is the field of (neural) controllable text generation (Prabhumoye et al., 2020). Here, the generation process of the language model is steered by another language model to e.g. adhere to a certain dialog structure, prevent toxic answers, or play a certain persona. Keskar et al. (2019) use control codes added to the prompt to steer the generation. Dathathri et al. (2020) propose the use of an external supervised classifier to control the generation. Chan et al. (2021) also utilize an external classifier but trains in a self-supervised setting. (Ghazvininejad et al., 2017; Holtzman et al., 2018) re-weight the probability distributions at each step of generating the output sequence. For further work in this field see the surveys by Prabhumoye et al. (2020); Zhang et al. (2023).
@@ 6 Experiments
We performed experiments on three free-form question-answering datasets, covering a broad range of question-answering settings. To be concrete, we use the over 800 closed-book questions in TruthfulQA (Lin et al., 2022a) corresponding to whole sentence answers, the almost 8,000 open-book questions in the development split of CoQA (Reddy et al., 2019) corresponding to medium to shorter length answers, and about 8,000 closed-book questions in the training split of TriviaQA (Joshi et al., 2017) corresponding to short, precise answers. We use a 5-shot, zero-shot, and 10-shot prompt for TruthfulQA, CoQA, and TriviaQA respectively. These datasets are frequently used as benchmarks for uncertainty estimation in NLG due to their strong correlation with human evaluations and the effective performance of “off-the-shelf” language models compared to tasks requiring fine-tuning (Kuhn et al., 2023; Goyal et al., 2023). Each of the three datasets was evaluated with the OPT model family (Zhang et al., 2022), with model sizes ranging from 2.7 to 30 billion parameters. Related work suggests that performance trends generalize across transformer-based model families (Duan et al., 2023; Manakul et al., 2023). In general, the four language models and three datasets assess the performance of uncertainty estimation methods in NLG across varying model sizes, output sequence lengths, and both open-book and closed-book settings.
@@
Following Kuhn et al. (2023); Lin et al. (2023); Duan et al. (2023), the quality of an uncertainty estimator is evaluated by how well it correlates with the respective correctness of an answer of the language model; correct answers should be assigned a lower uncertainty estimator than incorrect answers. Thus, the correctness of the initial output sequence $\bm{y}^{\prime}$ has to be evaluated. We follow the evaluation protocol of current work and utilize the statistics-based metrics Rouge-L and Rouge-1 (Lin, 2004) as well as the transfer learning-based metric BLEURT (Sellam et al., 2020), each with ten different correctness thresholds ranging from $0.1$ to $1.0$ (exact match). In general, the correctness of $\bm{y}^{\prime}$ is computed as [max score to a true reference answer] - [max score to a false reference answer]. While TurthfulQA provides both true and false reference answers, CoQA and TriviaQA only provide true reference answers. The initial output sequence $\bm{y}^{\prime}$ is sampled using a beam search with five beams. It also serves as the first output sequence utilized by every uncertainty estimation method. The Area Under Receiver Operating Characteristic (AUROC) is used as a metric for classifying the correct vs. incorrect answers, using the respective uncertainty estimate as a score. The higher the AUROC, the higher the correlation between the uncertainty estimate and the correctness of the answers.
@@
We compare our method against two streams of current predictive uncertainty estimation in NLG. First, methods that directly utilize the predictive entropy of the output distribution on a token level. These are Predictive Entropy (PE), Length-Normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2021), and Shifting Attention to Relevance (SAR) (Duan et al., 2023). Second, methods that utilize semantic entropy on a sequence level. Thereby, output sequences are generated with multinomial sampling (SEMS) (Kuhn et al., 2023) or with diverse beam search (SEDBS) (Vijayakumar et al., 2018). Although diverse beam search has not explicitly been proposed for uncertainty estimation in NLG, we added it as a more traditional sampling method that enforces diversity among output sequences. For each dataset, we performed a hyperparameter search on SEMS temperature $\in\{0.25,0.5,1.0,1.5,2.0\}$ and SEDBS penalty term $\in\{0.2,0.5,1.0\}$.
@@
Unlike current methods that rely on finding the optimal sampling temperature or penalty term, SDLG does not require hyperparameter tuning as it controls the sampling. One has only to decide on the ranking method for the three individual token scores. We empirically found that the performance of SDLG is quite robust with respect to the weighting of the token scores. Therefore, throughout the experiments, we derive the final token score ranking by straightforwardly averaging the three individual token scores. To compute the token scores for semantic diversity as discussed in Sec. 4, we utilize the same NLI model DeBERTa (Williams et al., 2018; He et al., 2021) that is also used for predicting semantic equivalences and determining semantic clusters. We also note that substituting initial tokens not corresponding to the beginning of a word is often impractical. Consequently, we exclusively apply substitutions to tokens at the beginning of a word.
Table 1: AUROC using different uncertainty measures as a score to distinguish between correct and incorrect answers. SEMS uses the improper semantic entropy estimator implemented by Kuhn et al. (2023) while ${}_{\ \cdots}^{(*)}$ use the proper semantic entropy estimator as we introduced in Sec. 3, with different sampling strategies. The threshold of the correctness metric Rouge-L (F1 score) is set to $0.5$. Each method uses ten output sequences for assigning an uncertainty estimate.
Dataset Model LN-PE PE SAR SEMS SE()${}_{MS}^{(*)}$ SE()${}_{DBS}^{(*)}$ SE()${}_{\texttt{SDLG}{}}^{(*)}$
TruthfulQA OPT-2.7b .439 .517 .611 .405 .846 .686 .920
OPT-6.7b .446 .510 .555 .512 .781 .637 .881
OPT-13b .676 .712 .775 .453 .896 .819 .956
OPT-30b .482 .542 .517 .438 .864 .788 .927
CoQA OPT-2.7b .717 .693 .733 .717 .744 .697 .744
OPT-6.7b .728 .703 .748 .739 .764 .714 .759
OPT-13b .723 .697 .747 .743 .758 .720 .760
OPT-30b .732 .698 .742 .745 .767 .713 .768
TriviaQA OPT-2.7b .769 .787 .785 .781 .804 .808 .809
OPT-6.7b .790 .805 .804 .803 .822 .823 .829
OPT-13b .807 .820 .819 .824 .838 .841 .845
OPT-30b .799 .812 .815 .817 .831 .837 .840
@@
Our method largely outperforms all current methods when utilizing each correctness metric Rouge-L (F1 score), Rouge-1 (F1 score), and BLEURT. The performance improvements of our method persist across different correctness thresholds as well as the number of samples considered. It can be observed that simple token-level diversity enforced by higher temperatures in multinomial sampling (Kuhn et al., 2023) or by diverse beam search (Vijayakumar et al., 2018) is insufficient for capturing semantic diversity essential for uncertainty estimation in NLG. The results summarized in Tab. 1 also show that using the proper semantic entropy estimator for the semantic entropy (SE()${}_{MS}^{(*)}$) outperforms the improper estimator (SEMS) when generating output sequences via multinomial sampling, as outlined in Sec. 3.
@@
Our method results in a $19\%$ increase of semantic clusters after generating the second output sequence, as well as a $74\%$ increase of semantic clusters after the tenth output sequence, compared to multinomial sampling with the highest-performing temperature used by the current state-of-the-art method (Kuhn et al., 2023), when averaging across all CoQA instances. This is because SDLG explicitly searches for output sequences with different semantic meanings and practically does not sample the same output sequence twice.
@@
Our experiments show that SDLG is sample efficient. Using a single output sequence generated with SDLG usually already performs better than using up to ten output sequences generated with other methods. While it is true that generating additional output sequences with SDLG requires extra computation for the NLI model to obtain the token score ranking, the primary computational effort lies in the generation of output sequences. The computational expenses of a single forward and backward pass through the NLI model with a few hundred million parameters is minor compared to generating only a single token using a language model with billions of parameters. Since our method deterministically changes a specific token within the initial output sequence, preceding tokens do not have to be regenerated again, but only subsequent ones. This results in SDLG requiring at least an average of $15\%$ (TruthfulQA), $33\%$ (CoQA), and $21\%$ (TriviaQA) fewer flops compared to multinomial sampling that is used by the current state-of-the-art method (Kuhn et al., 2023). The method SAR (Duan et al., 2023) even adds computational cost to multinomial sampling by utilizing the NLI model to evaluate the relevance of each token in the initial output sequence through a “leave-one-out” approach. In general, it can be observed that the advantage of our method over the current methods further increases with longer output sequences and larger language model sizes.
Experimental details and insights into the two ablation studies on semantic clusters and computational expenses can be found in Sec. D in the appendix.
@@ 7 Conclusion
We introduce SDLG to improve uncertainty estimation in NLG. SDLG substitutes tokens in the initial output sequence that are likely to lead to a change in semantic meaning. Unlike standard multinomial sampling, this targeted approach effectively samples likely output sequences with different semantic meanings, capturing important information for estimating semantic uncertainty. Our experiments on free-form question-answering datasets demonstrate that SDLG not only increases the overall quality of the uncertainty estimator but is also computationally more efficient.
Future work should investigate the performance of SDLG on NLG tasks with longer output sequences, such as summarization. However, evaluating the correctness of longer output sequences is a challenging problem in itself. Furthermore, the assumption that each sentence can be attributed to a single semantic cluster might be overly strict. Using continuous sentence similarities could allow for a relaxed semantic cluster assignment, which would only require minimal changes to our method. Moreover, the semantic difference is currently only conditioned on the first output sequence. We expect to see further improvements when taking into account the semantics of all previously generated output sequences as well. Lastly, this work focuses on estimating the aleatoric semantic uncertainty. Future work should investigate how to effectively assess the epistemic semantic uncertainty (see Eq. (3)).
Despite these remaining challenges, SDLG already offers notable advantages over prior methods. First, SDLG controls the sampling instead of depending on chance to obtain diverse samples. It eliminates the necessity of searching for an optimal sampling temperature, a hyperparameter important for the performance of all prior methods. Second, the advantageous sampling reduces the required number of samples for high-quality uncertainty estimation, while also being computationally efficient. Overall, SDLG has the potential to considerably enhance the applicability of uncertainty estimation in NLG.
@@ Acknowledgements
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation. Kajetan Schweighofer acknowledges travel support from ELISE (GA no 951847).

### v2

@@ 5 Related Work
Several works utilized the language model itself to obtain a prediction of their uncertainty, whether that be numerical or verbal (Mielke et al., 2022; Lin et al., 2022b; Kadavath et al., 2022; Cohen et al., 2023a; Ganguli et al., 2023; Ren et al., 2023; Tian et al., 2023). Cohen et al. (2023b) utilize cross-examination, where one language model generates the output sequence, and the other language model acts as an examiner to assess the uncertainty. Zhou et al. (2023) investigate the behavior of language models when expressing their (un)certainty.
A large body of work focuses on sampling a set of output sequences to obtain sampling-based uncertainty estimators. Xiao and Wang (2021); Malinin and Gales (2021); Hou et al. (2023) incorporate both aleatoric and epistemic estimates of uncertainty, where epistemic uncertainty due to model selection is considered. While Kuhn et al. (2023); Duan et al. (2023); Farquhar et al. (2024) evaluate the aleatoric uncertainty only under a single given language model, they take the semantic equivalence of potential output sequences into account. Manakul et al. (2023) also sample a set of output sequences but utilize them as input to another language model to assess the uncertainty. Another approach to uncertainty estimation in NLG is conformal prediction (Quach et al., 2023), where a stopping rule for generating output sequences is calibrated. Additionally, Xiao et al. (2022) empirically analyze how factors such as model architecture and training details influence the uncertainty estimates in language models.
Complementary methods.
Related work on uncertainty estimation in NLG aims to propose improved uncertainty measures compared to semantic entropy and could benefit from more diverse samples. For instance, Lin et al. (2023) propose leveraging the similarity between output sequences, while Chen et al. (2024) introduce EigenScore, which utilizes embeddings of output sequences. These approaches could integrate SDLG to generate the output sequences needed to compute their uncertainty estimators, making them complementary to SDLG.
@@
Li et al. (2016) propose an alternative training procedure of language models to avoid generic, input-independent output sequences and increase diversity. Diverse beam search (Vijayakumar et al., 2018) optimizes for a diversity-augmented objective across beam groups, based on diversity heuristics. Ippolito et al. (2019) compare diversity-encouraging decoding strategies. Nucleus sampling (Holtzman et al., 2020) generates higher quality as well as more diverse output sequences but does not explicitly encourage semantic diversity. Contrastive decoding (Li et al., 2023) utilizes a second, weaker language model, where the decoding algorithm favors tokens generated by the stronger model and penalizes tokens generated by the weaker model. Tam (2020) utilize semantic clustering during beam search, which is used to prune beams and diversify the remaining candidates. However, this only indirectly steers towards more diversity and relies on the diversity of the initial beams.
Closely related, but not directly targeting the semantic diversity of output sequences, is the field of (neural) controllable text generation (Prabhumoye et al., 2020). Here, the generation process of the language model is steered by another language model to e.g. adhere to a certain dialog structure, prevent toxic answers, or play a certain persona. Keskar et al. (2019) use control codes added to the prompt to steer the generation. Dathathri et al. (2020) propose the use of an external supervised classifier to control the generation. Chan et al. (2021) also utilize an external classifier but train in a self-supervised setting. (Ghazvininejad et al., 2017; Holtzman et al., 2018) re-weight the probability distributions at each step of generating the output sequence. For further work in this field, see the surveys by Prabhumoye et al. (2020); Zhang et al. (2023).
[…]
TruthfulQA 2.7b .439 .517 .611 .405 .846 .686 .920
6.7b .446 .510 .555 .512 .781 .637 .881
13b .676 .712 .775 .453 .896 .819 .956
30b .482 .542 .517 .438 .864 .788 .927
CoQA 2.7b .717 .693 .733 .717 .744 .697 .744
6.7b .728 .703 .748 .739 .764 .714 .759
13b .723 .697 .747 .743 .758 .720 .760
30b .732 .698 .742 .745 .767 .713 .768
TriviaQA 2.7b .769 .787 .785 .781 .804 .808 .809
6.7b .790 .805 .804 .803 .822 .823 .829
13b .807 .820 .819 .824 .838 .841 .845
30b .799 .812 .815 .817 .831 .837 .840
@@ 6 Experiments
To ensure a fair comparison of different uncertainty estimation methods within the scope of our computational budget, we decided to align the evaluation tasks with current work by focusing on free-form question answering. Thus, we performed experiments on three datasets that cover a broad range of question answering settings. To be concrete, we use the over 800 closed-book questions in TruthfulQA (Lin et al., 2022a) corresponding to whole sentence answers, the almost 8,000 open-book questions in the development split of CoQA (Reddy et al., 2019) corresponding to medium to shorter length answers, and about 8,000 closed-book questions in the training split of TriviaQA (Joshi et al., 2017) corresponding to short, precise answers. We use a 5-shot, zero-shot, and 10-shot prompt for TruthfulQA, CoQA, and TriviaQA, respectively. These datasets are frequently used as benchmarks for uncertainty estimation in NLG due to their strong correlation with human evaluations and the effective performance of “off-the-shelf” language models compared to tasks requiring fine-tuning (Kuhn et al., 2023; Goyal et al., 2023). Each of the three datasets was evaluated with the OPT model family (Zhang et al., 2022), with model sizes ranging from 2.7 to 30 billion parameters. Related work suggests that performance trends generalize across transformer-based model families (Duan et al., 2023; Manakul et al., 2023). In general, the four language models and three datasets assess the performance of uncertainty estimation methods in NLG across varying model sizes, output sequence lengths, and both open-book and closed-book settings.
@@
Effective uncertainty measures should accurately reflect the reliability of answers generated by the language model. Higher uncertainty more likely leads to incorrect generations. Thus, to evaluate the performance of an uncertainty estimator, we assess how well it correlates with the correctness of the language model’s answers; correct answers should be assigned a lower uncertainty estimator than incorrect answers. To determine whether an answer is correct, it has to be compared to the respective ground truth answer. First, we utilize the statistics-based metrics Rouge-L and Rouge-1 (Lin, 2004), which take the initial output sequence and the ground truth answer to measure the longest common subsequence and the overlap of unigrams, respectively. Second, we utilize the learning-based metric BLEURT (Sellam et al., 2020) that uses a learned evaluation model to assess how well the initial output sequence conveys the meaning of the ground truth answer. It has to be noted that for each of these three metrics, correctness is computed as the difference between the maximum score assigned to a true reference answer and the maximum score assigned to a false reference answer. While TruthfulQA provides both true and false reference answers, CoQA and TriviaQA only provide true reference answers. The resulting scalar value reflects the correctness of the answer and is subsequently thresholded to yield a binary decision. To ensure a robust evaluation, we use a comprehensive set of ten correctness thresholds ranging from $0.1$ to $1.0$ (exact match). To measure the correlation between incorrectness of answers and the respective uncertainty estimates, we utilize AUROC, with higher values indicating better performance of the uncertainty estimator, as it reflects a stronger alignment between the correctness of the language model’s answers and their respective uncertainty estimates. Overall, this evaluation process follows established methodologies for assessing the performance of uncertainty measures in NLG (Kuhn et al., 2023; Lin et al., 2023; Duan et al., 2023; Bakman et al., 2024; Farquhar et al., 2024; Nikitin et al., 2024).
Figure 3: (a) AUROC using uncertainty measures across various numbers of samples as score to distinguish between correct and incorrect answers of the CoQA dataset. Solid and dotted lines indicate the performance when using the proper and improper semantic entropy estimator, respectively.
(b) Average number of semantic clusters found across various numbers of samples considered.
@@
We compare our method against methods that directly utilize the predictive entropy of the output distribution on a token level. These are Predictive Entropy (PE), Length-Normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2021), and Shifting Attention to Relevance (SAR) (Duan et al., 2023). We also compare against methods that utilize semantic entropy on a sequence level. Thereby, output sequences are generated with multinomial sampling (SEMS) (Kuhn et al., 2023) or with diverse beam search (SEDBS) (Vijayakumar et al., 2018). Although DBS has not explicitly been proposed for uncertainty estimation in NLG, we added it as a more traditional sampling method that enforces diversity among output sequences.
@@
Unlike current methods that rely on finding the optimal sampling temperature or penalty term, SDLG does not require hyperparameter tuning as it controls the sampling. One has only to decide on the ranking method for the three individual token scores. We empirically found that the performance of our method is quite robust with respect to the weighting of the token scores. Therefore, throughout the experiments, we derive the final token score ranking by straightforwardly averaging the three individual token scores. To compute the token scores for semantic diversity as discussed in Sec. 4, we utilize the same NLI model DeBERTa (Williams et al., 2018; He et al., 2021) that is also used for predicting semantic equivalences and determining semantic clusters.
@@
[…]
It can be observed that simple token-level diversity enforced by higher temperatures in MS (Kuhn et al., 2023) or by DBS (Vijayakumar et al., 2018) is insufficient for capturing semantic diversity essential for uncertainty estimation in NLG. However, since TriviaQA includes the shortest output sequences among the datasets, the advantages of our method are less pronounced. Although our method still outperforms the baseline methods on short output sequences, the true strengths of SDLG emerge in scenarios with longer output sequences, such as those in TruthfulQA, where simple token-level diversity is insufficient, and the ability of our method to explore semantic diversity on a sequence-level becomes critical. The smallest performance gap between SE with MS and SDLG occurs on the CoQA dataset. Therefore, we more closely examine the relationship between the performance and the sample size. We report MS with the optimal temperature ($t=0.5$) as well as at nearby temperatures for comparison. As shown in Fig. 3(a), while the performance difference is smaller when considering all ten generations, SDLG demonstrates strong performance even after sampling just one additional output sequence. This highlights the sample efficiency of our method.
Figure 4: Average number of Teraflops required for an increasing number of samples generated with SDLG vs. standard multinomial sampling (MS) and Shifting Attention to Relevance (SAR).
@@
Fig. 3(b) shows that our method results in a $19\%$ increase of semantic clusters after generating the second output sequence, as well as a $74\%$ increase of semantic clusters after the tenth output sequence, compared to multinomial sampling with the highest-performing temperature used by the current state-of-the-art method (Kuhn et al., 2023), when averaging across all CoQA instances. This is because SDLG explicitly searches for output sequences with different semantic meanings and practically does not sample the same output sequence twice.
@@
While it is true that generating additional output sequences with SDLG requires extra computation for the NLI model to obtain the token score ranking, the primary computational effort lies in the generation of output sequences. The computational expenses of a single forward and backward pass through the NLI model with a few hundred million parameters is minor compared to generating only a single token using a language model with billions of parameters. Since our method deterministically changes a specific token within the initial output sequence, preceding tokens do not have to be regenerated again, but only subsequent ones. This results in SDLG requiring at least an average of $15\%$ (TruthfulQA), $33\%$ (CoQA), and $21\%$ (TriviaQA) fewer flops compared to multinomial sampling that is used by the current state-of-the-art method (Kuhn et al., 2023). In general, the advantage of our method over the current methods further increases with longer output sequences and larger language model sizes, as illustrated in Fig. 4.
@@ 7 Conclusion
We introduce SDLG to improve uncertainty estimation in NLG. SDLG substitutes tokens in the initial output sequence that are likely to lead to a change in semantic meaning. Unlike standard multinomial sampling, this targeted approach effectively samples likely output sequences with different semantic meanings, capturing important information for estimating semantic uncertainty. Our experiments on free-form question answering datasets demonstrate that SDLG not only increases the overall quality of the uncertainty estimator but is also computationally more efficient.
Future work should address the assumption that each sentence can be attributed to a single semantic cluster, which may be overly restrictive. This represents an orthogonal issue that does not conflict with our method but can be addressed independently. Also, this work focuses on estimating the aleatoric semantic uncertainty. Future work should investigate how to effectively assess epistemic semantic uncertainty, which is a challenging task in itself (see Eq. (3)).
Despite these remaining challenges, SDLG already offers notable advantages over prior methods. First, SDLG controls the sampling instead of depending on chance to obtain diverse samples. It eliminates the necessity of searching for an optimal sampling temperature, an important hyperparameter for all prior methods. Second, the advantageous sampling reduces the required number of samples for high-quality uncertainty estimation, while also being computationally most efficient per sample. Overall, SDLG considerably enhances the applicability of uncertainty estimation in NLG.
Ethics and Reproducibility Statement
We recognize the potential for language models to generate biased or harmful content if not properly managed. While SDLG aims to improve model reliability, we strongly advocate for its responsible use alongside bias mitigation strategies and content moderation techniques to ensure ethical and safe deployment.
To promote reproducibility, theoretical justifications are provided in Sec. 2, Sec. 3, and Sec. 4, with further mathematical derivations and proofs in the appendix. All datasets used are publicly available, and we utilize standard benchmarks to facilitate easy replication of our work. Upon publication, the source code for reproducing all experiments will be made publicly accessible.
@@ Acknowledgements
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. This research was funded in part by the Austrian Science Fund (FWF) [10.55776/COE12]. We thank the projects INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation. Kajetan Schweighofer acknowledges travel support from ELISE (GA no 951847).